OpenStudy (anonymous):

how do i find the solution set of x squared-4x-12=0

OpenStudy (anonymous):

Use the quadratic equation.

OpenStudy (anonymous):

which is?...

OpenStudy (anonymous):

\[ax^2 + bx + c = 0 \implies x = \frac{-b \pm \sqrt{b^2-4(a)(c)}}{2(a)}\]

OpenStudy (anonymous):

If you haven't learned that yet, you can solve this by factoring also.

OpenStudy (anonymous):

ok, can you help me with the factoring part please

OpenStudy (anonymous):

Oh, is it (x-6)(x+2) then x-6=0 add 6 to both sides = x=6 then x+2=0 sub. 2 from both sides and get x=-2 so it 6 and -2? \

OpenStudy (anonymous):

Yes. exactly!

OpenStudy (anonymous):

thankyouu !!

OpenStudy (anonymous):

Any time =)